A bicyclist starting at rest produces a constant angular acceleration of 1.50 rad/s2 for wheels that are 33.5 cm in radius.


(a) What is the bicycle's linear acceleration (in m/s2)? (Enter the magnitude.)

___ m/s2


(b) What is the angular speed of the wheels (in rad/s) when the bicyclist reaches 10.2 m/s?

___rad/s


(c) How many radians have the wheels turned through in that time?

___rad


(d) How far (in m) has the bicycle traveled?

___m


1
Expert's answer
2021-10-05T10:07:52-0400

"α=1.50 \;rad/s^2 \\

r = 33.5 \;cm"

(a)

"a = rα \\

= 33.5 \times 10^{-2} \times 1.50 \\

= 0.5025 \;m/s^2"

(b)

"v=rω \\

ω = \frac{v}{r} \\

= \frac{10.2}{0.335} \\

= 30.44 \;rad/s"

(c)

"v=u+at \\

u=0 \;(at \;rest) \\

10.2 = 0 + 0.5025t \\

t = 20.30 \;sec \\

ΔQ=\frac{1}{2} αt^2 \\

= \frac{1}{2} \times 1.5 \times (20.30)^2 \\

= 309.06 \;rad"

(d) Distance

"s=rΔQ \\

= 0.335 \times 309.06 \\

= 103.53 \;m"